# Pixel-Wise Recognition for Holistic Surgical Scene Understanding

### 20 Jan 2024 v1

@@ 2 Related Work
2.1 Surgical Workflow Analysis
Over the years, the evolution of datasets for surgical workflow analysis has substantially impacted the field. In 2014, the JIGSAWS dataset (Gao et al. (2014)) emerged as a pioneering effort by combining video and kinematic data with annotations for three fundamental surgical tasks in robotic surgery and fifteen atomic surgical gestures. Nevertheless, this dataset was acquired in a simulated training laboratory, distancing it from an actual surgical setting. Responding to the increasing demand for in vivo human surgical data, the Cholec80 dataset (Twinanda et al. (2016)) introduced a collection of laparoscopic cholecystectomy surgeries with annotations for surgical phase and tool presence recognition. Simultaneously, TUM LapChole (Stauder et al. (2017)), another dataset for laparoscopic cholecystectomies, also labeled surgical phases within a smaller collection of surgeries.
Notably, Cholec80 and TUM LapChole datasets focus on traditional laparoscopic procedures, distinguishing them from robot-assisted interventions. The Nephrec9 dataset (Nakawala (2017)) was introduced as the first dataset for robot-assisted interventions, specifically Robot-Assisted Partial Nephrectomy (RAPN) procedures. Remarkably, the Nephrec9 dataset provides temporal annotations at a more granular level, offering information about surgical steps instead of broader surgical phases.
@@ 2.1 Surgical Workflow Analysis
Following this effort, the HeiChole dataset (Wagner et al. (2023)) presented a detailed collection of 33 laparoscopic cholecystectomies and is also aimed at multi-level understanding by presenting phase recognition along with frame-level instrument and action presence recognition. While more specialized in its scope than the HeiCo dataset, HeiChole refined the understanding of surgical phases, along with the Cholec80 dataset (Twinanda et al. (2016)). Still, HeiChole did not include crucial spatially localized information or step recognition within surgical procedures.
@@
Even though the previously mentioned datasets have offered valuable insights into surgical workflow analysis within their specific surgical domains, they all study a single long-term task, and many of them confine their video-level annotations to broader procedural segments, such as surgical phases or steps. The MISAW dataset (Huaulmé et al. (2021)) emerged as a crucial addition to surgical workflow analysis datasets to address these gaps in existing datasets. Focused on micro-surgical anastomosis on artificial blood vessels, MISAW encompassed 27 sequences with videos, kinematic data, and temporal annotations across three granularity levels: phases, steps, and specific activities that are constituted by a surgical gesture, a target, and a surgical instrument. This dataset marked a significant stride towards enabling multi-level surgical workflow analysis. However, the artificial nature of the data limits its applicability to real-time surgical scene understanding.
@@
In contrast, the RARP-45 dataset (van Amsterdam et al. (2022)) stands out as a significant contribution by addressing the need for finer-grained temporal annotation within human in vivo surgical domains, specifically in robot-assisted interventions. This dataset comprises a comprehensive collection of video and kinematics data derived from Robot-Assisted Radical Prostatectomies (RARP), enriched with temporal segment annotations for fine-grained bi-manual gesture recognition. Nevertheless, it solely offers surgical gesture annotations.
Within the same RAS domain, the recently introduced RARR dataset (Kolbinger et al. (2023)) addresses the need for higher-level understanding in RAS by introducing annotations for phase recognition in human in vivo robot-assisted rectal resection (RARR). It also allows for multi-level modeling with lower-level semantic segmentation annotations of anatomical target structures. Despite providing spatial and temporal annotations, it does not address instrument recognition or the multi-granularity of broken-down phases into steps or surgical actions.
@@
In exploring surgical workflow analysis advancements, we acknowledged the need for a single benchmark encompassing hierarchically various granularity levels of long- and short-term surgical workflow tasks. We especially recognize this need in datasets utilizing real RAS data on human subjects. Our dataset, as an extension of our previously proposed PSI-AVA dataset (Valderrama et al. (2022)), provides annotations with high-level temporal information regarding phases and their division into steps, a more detailed temporal understanding. Thus, ours is the only surgical workflow analysis dataset up to date that provides both levels of semantic understanding with human in vivo data. Additionally, our dataset contains detailed low-level spatio-temporal information on atomic actions, and, in this work, we include precise delineation of the different surgical instruments for a more detailed understanding of instrument poses and surgical gestures. Consequently, our work addresses multiple previous needs in surgical scene understanding by offering a holistic comprehension of the surgical workflow at different granularity levels on real RAS data on human subjects.
@@ 2.4 Surgical Scene Understanding with Vision Transformers
This success has extended to the domain of robot-assisted surgeries. Previous works have successfully included transformers in their approach. For instance, in the domain of instrument classification, Kondo (2021) proposed a technique employing a CNN for feature extraction along with a temporal module based on Transformers, offering insights into a temporal window of features. In a distinctive approach, Nwoye et al. (2022) introduced a Transformer-inspired method for classifying triplets, incorporating a Convolutional Neural Network (CNN) backbone and attention mechanisms for enhanced feature mapping. Later, Sharma et al. (2023a) expanded on this model by incorporating a temporal consistency module, which fused features within a temporal window. Recently, Sharma et al. (2023b) proposed a triplet detection model, leveraging Deformable DETR (Zhu et al. (2021)) and a ResNet50 (He et al. (2016)) coupled with a transformer encoder for feature extraction. The methodology included incorporating a graph to process these features and understand interactions and a Multi-Head Attention layer applied to the nodes’ features.
In phase recognition, Gao et al. (2021) utilized a CNN backbone for spatial feature extraction, coupled with temporal processing through a Temporal Convolutional Network (TCN) and refinement with Transformer layers. Similarly, Czempiel et al. (2021) integrated a CNN backbone and Transformer layers, introducing a regularization parameter to focus attention on high-confidence features. Additionally, Ding and Li (2022) explored phases as temporal segments, employing a combination of CNN, temporal convolutions, and Transformers to achieve an enhanced capture of semantics and interrelationship.
@@
However, these methodologies have not exploited the potential of transformers for a comprehensive understanding of surgical scenes, often relying on CNNs for feature extraction. Moreover, these approaches are designed for specific tasks rather than employing a general design that embraces the holistic nature of surgeries. Regarding these gaps, our previous works, Valderrama et al. (2022) and Ayobi et al. (2023), leveraged transformers for various tasks. This approach marked the introduction of architectures entirely based on Transformers, demonstrating satisfactory results in instrument detection, instrument segmentation, phase classification, and step classification tasks, achieving state-of-the-art performance. In this regard, our TAPIS model introduces a single method fully based on transformer architectures that can effectively tackle multiple surgical workflow analysis tasks within different levels of visual, temporal, and spatial understanding. Thus presenting a generalized state-of-the-art transformer model for multi-granular surgical scene understanding.
@@ 3.2 Annotation Process
We classify phases and steps as long-term reasoning tasks, following the temporal breakdown outlined in Huaulmé et al. (2021). Phases represent the fundamental segments of surgical procedures, including key stages like preparation, organ manipulation, and anatomical reconstruction. Meanwhile, steps provide a more detailed breakdown within these phases, focusing on specific surgical activities and maneuvers, such as dissecting anatomical landmarks, making incisions, pulling sutures, and packing dissected structures. In the following Section, we explain the process of defining phases and step categories.
Conversely, our approach for localized atomic action recognition and surgical instrument segmentation focuses on short-term reasoning tasks. On the one hand, this facet includes identifying localized atomic actions within RARP procedures based on the atomic action definition of AVA (Gu et al. (2018)). Our medical experts stated a set of surgical atomic actions based on the most detailed and singular actions that could be performed in surgery by instruments such as pull, push, and hold, among others. Table 2 displays the final defined set of atomic action categories on the right side. On the other hand, surgical instrument segmentation involves precisely delineating and identifying surgical tools within the scene. We defined Surgical Instrument classes based on the da Vinci Robot user manual and catalog (Intuitive Inc. (2014, 2023)) and the instrument setup used during data collection (Section 3.1). We showcase the final set of surgical instrument categories in Table 2. By categorizing these aspects as short-term reasoning tasks, our dataset represents both the general and detailed facets of the RARP procedure, enabling a multi-level holistic understanding of surgical scenes in varying degrees of granularity. Due to the time-consuming nature of short-term task annotations and considering our resource constraints, we only annotated a subset of keyframes uniformly sampled through time with a fixed temporal stride, as explained in upcoming Sections 3.2.2 and 3.2.3.
@@ 3.2.1 Phase and Step Tasks Annotation
We then established the sequencing and hierarchy of phases based on standardized steps intrinsic to Robot-Assisted Radical Prostatectomy, a procedural protocol routinely practiced at Fundación Santa Fe de Bogotá. Subsequently, our surgical team reviewed the initial version of the phases and steps categories and hierarchies, involving scrutiny by two experienced urologists with specialized proficiency in Robotic-Assisted Surgery. This collaborative evaluation facilitated the refinement of the defined phases and steps, ensuring alignment with the intricate aspects of the surgical technique. The final set of phases and steps categories is portrayed in Table 2, and the hierarchical structure of phases and steps is visually represented in Figure A.13 in the Supplementary Material.
After establishing the labels and hierarchies of the long-term tasks, we conducted an annotation procedure that involved a team of two urologist residents specializing in transperitoneal radical prostatectomy surgeries for the five additional surgeries of our dataset. The annotation team independently annotated each surgical video, defining specific timestamps for the start and end times of phases and step intervals down to the hour, minute, and second, resulting in two annotations per video. In the case of the eight previously annotated surgeries, these were annotated by the team once more, resulting in two sets of annotations per surgery as well. When discrepancies arose between the annotations, the resident annotators collectively reviewed and reconciled differences to establish a consensus regarding the definitive delineation of phases and steps.
@@ 3.2.3 Atomic Action Recognition
The annotation process for atomic action recognition involved the team of four experts in surgical data science trained explicitly for this task with the guidance of the surgery experts team, following a detailed annotation protocol designed by domain experts. The annotation process consisted of three main stages: (1) Estimation of instrument location with bounding boxes and labeling of atomic actions, (2) Spatial refinement and validation by retrieving bounding boxes from the previously precisely annotated segmentation masks, and (3) Annotation consensus and rectification between independent annotators and guidance of experienced surgeons.
We underwent this process using the open-source Toronto Annotation Suite (Kar et al. (2021)) platform to annotate the same previously selected frames for surgical instrument segmentation. Since atomic action labeling explicitly requires temporal context, annotators defined atomic actions based on the 3-second clips centered on the frame of interest as initially done in AVA (Gu et al. (2018)). However, annotators could also access all of the surgical footage when necessary. Hence, the annotators first localized instruments with bounding boxes regardless of the instrument segmentation annotation process. After defining surgical instruments’ spatial location with bounding boxes, annotators identified and labeled combinations of maximum 3 atomic actions, treating them as discrete states that instruments could simultaneously undertake based on the predefined set of atomic action categories (Table 2).
@@
Additionally, we constructed theoretical co-occurrence matrices under the guidance of our experienced surgeons to establish valid presence relationships among the different atomic actions and between actions and instruments (Figure 23). These matrices ensured logical consistency in actions performed simultaneously by an instrument and verified correlations between specific actions executed by their annotated instrument type. We applied these matrixes to validate the annotated combinations of atomic actions, ensuring compatibility within sets of actions. In cases where conflicting atomic actions arose, or actions were incompatible with the corresponding instrument, the team collectively reviewed and deliberated to establish consensus for re-annotation.
@@ 3.3 Dataset Statistics
Figure 1 portrays an overview of our proposed tasks and annotations, and Table 2 presents our sets of semantic categories. The final ground truth annotations of GraSP comprise eleven (11) surgical phase categories, twenty-one (21) surgical steps, seven (7) instrument categories, and fourteen (14) atomic action categories. Therefore, our GraSP dataset proposes diverse classes for each task, integrating local and temporal visual understanding of surgeries. The hierarchical dendrogram of Figure A.13 of the Supplementary Material represents the relations between each phase and its nested steps, and it also includes the labels and names of all the categories (IDs) for each task in the GraSP dataset.
Figures 2 and 3 illustrate the distribution of the percentage of annotated frames (counted at 1fps) for each phase and step class, respectively. We also present transition matrices of the annotated categories and visualizations of the temporal progression of phases and steps in Figures B.15, B.16, B.17, and B.18 of the Supplementary Material. The frequency distribution of the long-term annotations for a particular class is highly influenced by the occurrence rate of a specific phase or step and the average duration of each occurrence. Notably, these distributions reveal a pronounced long-tail effect, indicating that some phase and step categories are significantly more prevalent than others. For instance, the Idle class is over-represented in the annotations of phases and steps compared to other categories. The transition matrices and temporal visualizations explain this trend, as they demonstrate that most phases and steps frequently transition from or to an Idle state. This tendency is typical and representative of Radical Prostatectomies, as surgeons pause between stages to change instruments, evacuate waste, clean and relocate the endoscope, or rest due to the complexity and duration of the procedures.
Fig. 2: Distribution of the percentage of frames per surgical phase category on each data split. We count the number of frames sampled at 1fps. We organize the graphic in descending order and present correspondence between phase IDs and phase labels in Figure A.13 of the Supplementary Material. The figure is best viewed in color.
Fig. 3: Distribution of the percentage of frames per surgical step category on each data split. We count the number of frames sampled at 1fps. We organize the graphic in descending order and present correspondence between step IDs and step labels in Figure A.13 of the Supplementary Material. The figure is best viewed in color.
@@
In addition, for Phase and Step annotations, we present the duration distribution of each continuous temporal segment for each phase and step category in seconds. We display the phase duration distribution in Figure 4 and the step duration distribution in Figure 5. These distributions and the visualizations in Figures B.17 and B.18 of the Supplementary Material underscore these tasks’ granularity and complexity differences. On the one hand, phases represent broader surgery segments, usually occurring once (though interrupted by Idle segments) and covering more general aspects of the procedure. Thus, they constitute a coarser level of understanding and exhibit longer average durations, with some continuous segments lasting over 20 minutes. On the other hand, steps recur in shorter segments repetitively, representing more specific parts of the surgery. Hence, step segments present shorter average durations, with the most extended segments spanning around 11 minutes, indicating a finer-grained level of detail.
@@
In this way, the observed statistical distributions of our long-term annotations, like the long-tail distribution in frame frequency and the significant inter- and intra-class variability in our phases and steps’ duration, present considerable challenges within our benchmark. Nevertheless, it is evident that these distributions accurately represent the realities of surgical scenarios, thus emphasizing the value of utilizing human in-vivo data sources and demonstrating the complexities involved in visually understanding surgical scenes.
@@
Furthermore, Figure 6 and Table B.18 (Supplementary Material) display the frequency distribution of the instrument instances per class, also demonstrating a natural long-tail effect. The reason for this distribution can also be observed in the co-occurrence matrices detailed in the Supplementary Material. The Monopolar Curved Scissors (MCS) and Bipolar Forceps (BF) are the most frequent instruments. This over-representation is primarily because these tools are the most actively used in robot-assisted surgeries, owing to their versatility and capacity to perform most surgical actions. They are also among the most crucial robotic instruments of the da Vinci Surgical System (dVSS). Following these, the Large Needle Driver (LND), another typical robotic instrument of the dVSS, also presents high frequencies as it is utilized for all suturing stages, but it has reduced functionalities compared to the MCS and BF. While also commonly used, the Prograsp Forceps (PF) and the Suction Instrument (SI) have more specific roles in surgery and often function passively and steadily. The latter is additional to the dVSS. Finally, instruments like the Clip Applier (CA) and other Laparoscopic Instruments (LI) are also additional to the dVSS and employed for specific purposes during certain stages of the surgery.
@@
Figure 7 illustrates the distribution of the number of instrument instances per video frame. We observe that each video frame contains 1 to 5 instrument instances (2.23 average of instrument instances per frame), with approximately 80% of frames containing between 2 and 3 instances. These statistics are characteristic of robotic surgeries, where surgeons mainly use two robotic instruments simultaneously with their hands. The additional presence of non-robotic instruments (SI, LI, and CA) manipulated by an assistant or a third robotic instrument left stationary increases the possible number of instances per frame and contributes to the diversity in surgical visual scenes. Situations involving four or five instruments are less common and usually indicate specific complex scenarios within the surgery. The upper limit of five instruments per frame is due to the inherent limitations of the dVSS and the specific complexities associated with Radical Prostatectomies.
@@
Finally, Figure 9 and Table B.19 (Supplementary Material) present the frequency and normalized distribution of the atomic action instances per class. Again, the distribution of atomic actions suggests a long-tail behavior where still (11), hold (5), and travel (13) are the most frequently occurring actions in the dataset. Indeed, during the surgery, most instruments are either used to hold and maintain tissues and objects in place or to move these tissues and objects to a different location. Later on, Push (9) and Pull (8) represent more specialized movements and thus are less present. Suction (12) is an action that can only be performed by the SI (as shown in the instrument-action co-occurrence matrix in Figure B.19 in the Supplementary Material). Nonetheless, when on the frame, the SI very frequently performs this action, which explains the higher frequency of this action. The remaining actions present lower frequencies as they are much more specialized and usually done by a specific type of instrument in shorter time frames. For instance, Cauterize (1) can only be performed by the BF and MCS, and Cut (3) is only performed by the MCS. Similarly, Open (6) can be done by all the instruments with tips (all except SI) but for very short durations. Lastly, the Other class (14) is the most scarce category since it is a composition of Staple and Wash, which can only be performed by the CA and the SI, respectively, and only happen in very particular moments of the surgery.
@@
Fig. 9: Distribution of the percentage of instances per atomic action category on each data split. We consider each annotated action within an instrument instance an independent atomic action instance. We organize the graphic in descending order and present the correspondence between atomic action IDs and atomic action labels in Figure A.13 of the Supplementary Material. The figure is best viewed in color.

### 26 Jan 2024 v2

@@ 2 Related Work
2.1 Surgical Workflow Analysis Benchmarks
Over the years, the evolution of datasets for surgical workflow analysis has substantially impacted the field. In 2014, the JIGSAWS dataset (Gao et al. (2014)) emerged as a pioneering effort by combining video and kinematic data with annotations for three fundamental surgical tasks in robotic surgery and fifteen surgical gestures. Nevertheless, this dataset was acquired in a simulated training laboratory, distancing it from an actual surgical setting. Responding to the demand for in vivo human surgical data, the Cholec80 dataset (Twinanda et al. (2016)) introduced a collection of laparoscopic cholecystectomy surgeries with annotations for surgical phase and tool presence recognition. Simultaneously, TUM LapChole (Stauder et al. (2017)), another dataset for laparoscopic cholecystectomies, also labeled surgical phases within a smaller collection of surgeries. Given that the previous datasets focus on traditional laparoscopic procedures, the Nephrec9 dataset (Nakawala (2017)) was introduced as the first surgical workflow analysis dataset with real RAS data of Robot-Assisted Partial Nephrectomy (RAPN) procedures. Remarkably, the Nephrec9 dataset provides temporal annotations at a more granular level, offering information about surgical steps instead of broader surgical phases.
@@ 2.1 Surgical Workflow Analysis Benchmarks
Following this effort, the HeiChole dataset (Wagner et al. (2023)) presented a detailed collection of 33 laparoscopic cholecystectomies and is also aimed at multi-level understanding by presenting phase recognition along with frame-level instrument and action presence recognition. While more specialized in its scope than the HeiCo dataset, HeiChole did not include crucial spatially localized information or step recognition within surgical procedures.
@@
Even though the previously mentioned datasets have offered valuable insights into surgical workflow analysis within their specific surgical domains, they all confine their video-level annotations to studying a single long-term task. The MISAW dataset (Huaulmé et al. (2021)) emerged as a crucial addition to surgical workflow analysis to address the previous gap in existing datasets. Focused on micro-surgical anastomosis on artificial blood vessels, MISAW encompassed 27 sequences with videos, kinematic data, and temporal annotations across three granularity levels: phases, steps, and specific activities that are constituted by a surgical gesture, a target, and a surgical instrument. This dataset marked a significant stride towards enabling multi-level surgical workflow analysis. However, the artificial nature of the data limits its applicability to real-time surgical scene understanding.
@@
In contrast, the RARP-45 dataset (van Amsterdam et al. (2022)) stands out as a significant contribution by addressing the need for finer-grained temporal annotation within human in vivo robot-assisted interventions. This dataset comprises a comprehensive collection of video and kinematics data derived from Robot-Assisted Radical Prostatectomies (RARP), enriched with temporal segment annotations for fine-grained bi-manual gesture recognition. Nevertheless, it solely offers surgical gesture annotations. Within the same RAS domain, the recently introduced RARR dataset (Kolbinger et al. (2023)) allows multi-level understanding by introducing annotations for higher-level phase recognition and lower-level anatomical targets segmentation in human in vivo robot-assisted rectal resection (RARR). Despite providing spatial and temporal annotations, it does not address instrument recognition or the multi-granularity of broken-down phases into steps or surgical actions.
@@
In exploring surgical workflow analysis advancements, we acknowledged the need for a single benchmark encompassing hierarchically various granularity levels of long- and short-term surgical workflow tasks. We especially recognize this need in datasets utilizing real RAS data on human subjects. Our dataset, as an extension of our previously proposed PSI-AVA dataset (Valderrama et al. (2022)), provides annotations with high-level temporal information regarding phases and their division into steps, a more detailed temporal understanding. Thus, ours is the only surgical workflow analysis dataset up to date that provides both levels of semantic understanding with human in vivo data. Additionally, our dataset contains detailed low-level spatio-temporal information on atomic actions, and, in this work, we include precise delineation of the different surgical instruments for a more detailed understanding of instrument poses and surgical gestures. Consequently, our work addresses multiple previous needs in surgical scene understanding by offering a holistic comprehension of the surgical workflow at different granularity levels with real RAS data on human subjects.
@@ 2.4 Surgical Scene Understanding with Vision Transformers
This success has extended to the domain of endoscopic surgeries since previous works have successfully included transformers in their approach. For instance, in the domain of instrument classification, Kondo (2021) proposed a technique employing a CNN for feature extraction along with a temporal module based on Transformers, offering insights into a temporal window of features. In a distinctive approach, Nwoye et al. (2022) introduced a Transformer-inspired method for classifying triplets, incorporating a Convolutional Neural Network (CNN) backbone and attention mechanisms for enhanced feature mapping. Later, Sharma et al. (2023a) expanded on this model by incorporating a temporal consistency module, which fused features within a temporal window. Recently, Sharma et al. (2023b) proposed a triplet detection model, leveraging Deformable DETR (Zhu et al. (2021)) and a ResNet50 (He et al. (2016)) coupled with a transformer encoder for feature extraction. The methodology included incorporating a graph to process these features and understand interactions and a Multi-Head Attention layer applied to the nodes’ features.
In phase recognition, Gao et al. (2021) utilized a CNN backbone for spatial feature extraction, coupled with temporal processing through a Temporal Convolutional Network (TCN) and refinement with Transformer layers. Similarly, Czempiel et al. (2021) integrated a CNN backbone and Transformer layers, introducing a regularization parameter to focus attention on high-confidence features. Additionally, Ding and Li (2022) explored phases as temporal segments, employing a combination of CNN, temporal convolutions, and Transformers to achieve an enhanced capture of semantics and interrelations.
@@
However, these methodologies have not exploited the potential of transformers for a comprehensive understanding of surgical scenes, often relying on CNNs for feature extraction. Moreover, these approaches are designed for specific tasks rather than employing a general design that embraces the holistic nature of surgeries. Regarding these gaps, our previous works, Valderrama et al. (2022) and Ayobi et al. (2023), leveraged transformers for various tasks. These methods marked the introduction of architectures entirely based on Transformers, demonstrating satisfactory results in instrument detection, instrument segmentation, phase classification, and step classification tasks, achieving state-of-the-art performance. In this regard, our TAPIS model introduces a single method fully based on transformer architectures that can effectively tackle multiple surgical workflow analysis tasks within different levels of visual, temporal, and spatial understanding. Thus presenting a generalized state-of-the-art transformer model for multi-granular surgical scene understanding.
@@ 3.2 Annotation Process
We classify phases and steps as long-term reasoning tasks, following the temporal breakdown outlined in Huaulmé et al. (2021). Phases represent the fundamental segments of surgical procedures, including key stages like preparation, organ manipulation, and anatomical reconstruction. Meanwhile, steps provide a more detailed breakdown within these phases, focusing on specific surgical activities and maneuvers, such as dissecting anatomical landmarks, making incisions, pulling sutures, and packing dissected structures. In the following subsection, we explain the process of defining phases and step categories.
Conversely, our approach for localized atomic action recognition and surgical instrument segmentation focuses on short-term reasoning tasks. On the one hand, this facet includes identifying localized atomic actions within RARP procedures based on the atomic action definition of AVA (Gu et al. (2018)). Our medical experts stated a set of surgical atomic actions based on the most detailed and singular actions that could be performed in surgery by instruments, such as pull, push, and hold, among others. Table 2 displays the final defined set of atomic action categories on the right side. On the other hand, surgical instrument segmentation involves precisely delineating and identifying surgical tools within the scene. We defined Surgical Instrument classes based on the da Vinci Robot user manual and catalog (Intuitive Inc. (2014, 2023)) and the instrument setup used during data collection (Section 3.1). We showcase the final set of surgical instrument categories in Table 2. By categorizing these aspects as short-term reasoning tasks, our dataset represents both the general and detailed facets of the RARP procedure, enabling a multi-level holistic understanding of surgical scenes in varying degrees of granularity. Due to the time-consuming nature of short-term task annotations and considering our resource constraints, we only annotated a subset of keyframes uniformly sampled through time with a fixed temporal stride, as explained in upcoming Subsections 3.2.2 and 3.2.3.
@@ 3.2.1 Phase and Step Tasks Annotation
We then established the sequencing and hierarchy of phases based on standardized steps intrinsic to Robot-Assisted Radical Prostatectomy, a procedural protocol routinely practiced at Fundación Santa Fé de Bogotá. Subsequently, our surgical team reviewed the initial version of the phases and steps categories and hierarchies, involving scrutiny by two experienced urologists with specialized proficiency in Robotic-Assisted Surgery. This collaborative evaluation facilitated the refinement of the defined phases and steps, ensuring alignment with the intricate aspects of the surgical technique. The final set of phases and steps categories is portrayed in Table 2, and the hierarchical structure of phases and steps is visually represented in Figure 2 in the Supplementary Material.
After establishing the labels and hierarchies of the long-term tasks, we conducted an annotation procedure that involved a team of two urologist residents specializing in transperitoneal radical prostatectomy surgeries. The annotation team independently annotated each surgical video, defining specific timestamps for the start and end times of phases and step intervals down to the hour, minute, and second, resulting in two annotations per video. In the case of the eight previously annotated surgeries, these were annotated by the team once more, resulting in two sets of annotations per surgery as well. When discrepancies arose between the annotations, the resident annotators collectively reviewed and reconciled differences to establish a consensus regarding the definitive delineation of phases and steps.
@@ 3.2.3 Atomic Action Recognition
The annotation process for atomic action recognition involved the team of four experts in surgical data science trained explicitly for this task with the guidance of the surgeons team, following a detailed annotation protocol designed by domain experts. The annotation process consisted of three main stages: (1) Estimation of instrument location with bounding boxes and labeling of atomic actions, (2) Spatial refinement and validation by retrieving bounding boxes from the previously precisely annotated segmentation masks, and (3) Annotation consensus and rectification between independent annotators and guidance of experienced surgeons.
We underwent this process using the open-source Toronto Annotation Suite (Kar et al. (2021)) platform to annotate the same previously selected frames for surgical instrument segmentation. Since atomic action labeling explicitly requires temporal context, annotators defined atomic actions based on the 3-second clips centered on the frame of interest as initially done in AVA (Gu et al. (2018)). However, annotators could also access all of the surgical footage when necessary. Hence, the annotators first localized instruments with bounding boxes regardless of the instrument segmentation annotation process. After defining surgical instruments’ spatial locations with bounding boxes, annotators identified and labeled combinations of maximum 3 atomic actions, treating them as discrete states that instruments could simultaneously undertake based on the predefined set of atomic action categories (Table 2).
@@
Additionally, we constructed theoretical co-occurrence matrices under the guidance of our experienced surgeons to establish valid presence relationships among the different atomic actions and between actions and instruments (Figure 11 of the Supplementary Material). These matrices ensured logical consistency in actions performed simultaneously by an instrument and verified correlations between specific actions executed by their annotated instrument type. We applied these matrixes to validate the annotated combinations of atomic actions, ensuring compatibility within sets of actions. In cases where conflicting atomic actions arose, or actions were incompatible with the corresponding instrument, the team collectively reviewed and deliberated to establish consensus for re-annotation.
@@ 3.3 Dataset Statistics
Figure 1 portrays an overview of our proposed tasks and annotations, and Table 2 presents our sets of semantic categories. The final ground truth annotations of GraSP comprise eleven (11) surgical phase categories, twenty-one (21) surgical steps, seven (7) instrument categories, and fourteen (14) atomic action categories. Therefore, our GraSP dataset proposes diverse classes for each task, integrating local and temporal visual understanding of surgeries. The hierarchical dendrogram of Figure 2 of the Supplementary Material represents the relations between each phase and its nested steps, and it also includes the labels and names of all the categories (IDs) for each task in the GraSP dataset.
Figures 2 and 3 illustrate the distribution of the percentage of annotated frames (counted at 1fps) for each phase and step class, respectively. We also present transition matrices of the annotated categories and visualizations of the temporal progression of phases and steps in Figures 5, 6, 7, and 9 of the Supplementary Material. The frequency distribution of the long-term annotations for a particular class is highly influenced by the occurrence rate of a specific phase or step and the average duration of each occurrence. Notably, these distributions reveal a pronounced long-tail effect, indicating that some phase and step categories are significantly more prevalent than others. For instance, the Idle class is over-represented in the annotations of phases and steps compared to other categories. The transition matrices and temporal visualizations explain this trend, as they demonstrate that most phases and steps frequently transition from or to an Idle state. This tendency is typical and representative of Radical Prostatectomies, as surgeons pause between stages to change instruments, evacuate waste, clean and relocate the endoscope, or rest due to the complexity and duration of the procedures.
Fig. 2: Distribution of the percentage of frames per surgical phase category on each data split. We count the number of frames sampled at 1fps. We organize the graphic in descending order and present correspondence between phase IDs and phase labels in Table 2. The figure is best viewed in color.
Fig. 3: Distribution of the percentage of frames per surgical step category on each data split. We count the number of frames sampled at 1fps. We organize the graphic in descending order and present correspondence between step IDs and step labels in Table 2. The figure is best viewed in color.
@@
In addition, for Phase and Step annotations, we present the duration distribution of each continuous temporal segment for each phase and step category in seconds. We display the phase duration distribution in Figure 4 and the step duration distribution in Figure 5. These distributions and the visualizations in Figures 7 and 9 of the Supplementary Material underscore these tasks’ granularity and complexity differences. On the one hand, phases represent broader surgery segments, usually occurring once (though interrupted by Idle segments) and covering more general aspects of the procedure. Thus, they constitute a coarser level of understanding and exhibit longer average durations, with some continuous segments lasting over 20 minutes. On the other hand, steps recur in shorter segments repetitively, representing more specific parts of the surgery. Hence, step segments present shorter average durations, with the most extended segments spanning around 11 minutes, indicating a finer-grained level of detail.
@@
In this way, the observed statistical distributions of our long-term annotations, like the long-tail distribution in frame frequency and the significant inter- and intra-class variability in our phases and steps’ duration, present considerable challenges within our benchmark. Nevertheless, it is evident that these distributions accurately represent the realities of surgical scenarios, thus emphasizing the value of utilizing human in vivo data sources and demonstrating the complexities involved in visually understanding surgical scenes.
@@
Furthermore, Figure 6 and Table 3 (Supplementary Material) display the frequency distribution of the instrument instances per class, also demonstrating a natural long-tail effect. The reason for this distribution can also be observed in the co-occurrence matrices detailed in the Supplementary Material. The Monopolar Curved Scissors (MCS) and Bipolar Forceps (BF) are the most frequent instruments. This over-representation is primarily because these tools are the most actively used in robot-assisted surgeries, owing to their versatility and capacity to perform most surgical actions. They are also among the most crucial robotic instruments of the da Vinci Surgical System (dVSS). Following these, the Large Needle Driver (LND), another typical robotic instrument of the dVSS, also presents high frequencies as it is utilized for all suturing stages, but it has reduced functionalities compared to the MCS and BF. While also commonly used, the Prograsp Forceps (PF) and the Suction Instrument (SI) have more specific roles in surgery and often function passively and steadily. The latter is additional to the dVSS. Finally, instruments like the Clip Applier (CA) and other Laparoscopic Instruments (LI) are also additional to the dVSS and employed for specific purposes during certain stages of the surgery.
Figure 7 and Table 4 (Supplementary Material) illustrate the distribution of the number of instrument instances per video frame. We observe that each video frame contains 1 to 5 instrument instances (2.23 average of instrument instances per frame), with approximately 80% of frames containing between 2 and 3 instances. These statistics are characteristic of robotic surgeries, where surgeons mainly use two robotic instruments simultaneously with their hands. The additional presence of non-robotic instruments (SI, LI, and CA) manipulated by an assistant or a third robotic instrument left stationary increases the possible number of instances per frame and contributes to the diversity in surgical visual scenes. Situations involving four or five instruments are less common and usually indicate specific complex scenarios within the surgery. The upper limit of five instruments per frame is due to the inherent limitations of the dVSS and the specific complexities associated with Radical Prostatectomies.
@@
Finally, Figure 9 and Table 5 (Supplementary Material) present the frequency and normalized distribution of the atomic action instances per class. Again, the distribution of atomic actions suggests a long-tail behavior where still (11), hold (5), and travel (13) are the most frequently occurring actions in the dataset. Indeed, during the surgery, most instruments are either used to hold and maintain tissues and objects in place or to move these tissues and objects to a different location. Later on, Push (9) and Pull (8) represent more specialized movements and thus are less present. Suction (12) is an action that can only be performed by the SI (as shown in the instrument-action co-occurrence matrix in Figure 14 in the Supplementary Material). Nonetheless, when on the frame, the SI very frequently performs this action, which explains the higher frequency of this action. The remaining actions present lower frequencies as they are much more specialized and usually done by a specific type of instrument in shorter time frames. For instance, Cauterize (1) can only be performed by the BF and MCS, and Cut (3) is only performed by the MCS. Similarly, Open (6) can be done by all the instruments with tips (all except SI) but for very short durations. Lastly, the Other class (14) is the most scarce category since it is a composition of Staple and Wash, which can only be performed by the CA and the SI, respectively, and only happen in very particular moments of the surgery.
@@
Fig. 9: Distribution of the percentage of instances per atomic action category on each data split. We consider each annotated action within an instrument instance an independent atomic action instance. We organize the graphic in descending order and present the correspondence between atomic action IDs and atomic action labels in Table 2. The figure is best viewed in color.